OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 64 cubic centimeters and 125 cubic centimeters?

OpenStudy (anonymous):

1:5 2:3 4:5 16:25

OpenStudy (mertsj):

Since the similarity factor was cubed in order to find the volume, take the cube root of the volume ratio to find the similarity ratio.

OpenStudy (anonymous):

is it the first one

OpenStudy (mertsj):

\[\sqrt[3]{\frac{64}{125}}=\frac{4}{5}\]

OpenStudy (anonymous):

i see thank u

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

can u help me with this one

OpenStudy (anonymous):

.A rectangular prism is truncated removing 20% of its height. The original height of the prism was 100 inches and it has a length of 5 inches and a width of 10 inches. What is the volume of the truncated prism? Answer 2,000 in3 3,000 in3 4,000 in3 5,000 in3